# latexText page

### v1

@@ 5.2 Model evaluation
(1) ES contribution (AMZN)
HS 33.603 6 0.179 0.910 0.090 yellow
GARCH.BU 18.965 3 0.810 0.595 0.405 yellow
GARCH.TD 20.026 4 0.698 0.651 0.349 yellow
HAR.GARCH 13.351 1 0.579 0.289 0.711 yellow
CR.LSE 23.075 5 0.224 0.888 0.112 yellow
CR.OPT 17.438 2 —- —- —- —-
(2) ES contribution (GOOGL)
HS 28.672 6 0.067 0.966 0.034 red
GARCH.BU 13.196 4 0.941 0.529 0.471 yellow
GARCH.TD 12.117 2 0.655 0.328 0.672 yellow
HAR.GARCH 15.629 5 0.364 0.818 0.182 yellow
CR.LSE 11.091 1 0.440 0.220 0.780 yellow
CR.OPT 13.005 3 —- —- —- —-
(3) ES contribution (TSLA)
HS 126.941 6 0.018 0.991 0.009 red
GARCH.BU 83.052 4 0.035 0.982 0.018 red
GARCH.TD 64.930 3 0.377 0.812 0.188 yellow
HAR.GARCH 95.799 5 0.137 0.932 0.068 yellow
CR.LSE 61.090 2 0.490 0.755 0.245 yellow
CR.OPT 54.984 1 —- —- —- —-
We next describe our traditional backtests. Denote by $\left(\widehat{ESC}_{1,t},\dots,\widehat{ESC}_{d,t},\widehat{VaR}_{t},\widehat{ES}_{t}\right)$, $t\in\mathbb{T}_{\text{out}}$, the tuple of forecasted risk functionals. For total VaR, let $\operatorname{V}_{t}^{\text{VaR}}=\operatorname{V}^{\text{VaR}}\left(\widehat{VaR}_{t},S_{t}\right)$, where $\operatorname{V}^{\text{VaR}}$ is as defined in Proposition 2. We construct the DM-type test statistic $\sqrt{T}\bar{\operatorname{V}}_{T}^{\text{VaR}}/\hat{\sigma}_{T}^{\text{VaR}}$, where $\bar{\operatorname{V}}_{T}^{\text{VaR}}=(1/T)\sum_{t=1}^{T}\operatorname{V}_{t}^{\text{VaR}}$, and $\hat{\sigma}_{T}^{\text{VaR}}$ is the HAC estimator of the standard deviation of $\sqrt{T}\bar{\operatorname{V}}_{T}^{\text{VaR}}$ as considered in the comparative backtests above. Following Section 2.2 of Nolde and Ziegel (2017), we use a normal distribution as a null distribution to test unconditional calibration, sub-calibration, and super-calibration, which are associated with precise estimation, over-estimation, and under-estimation, respectively, for this identification function of VaR; see Section 2.3. Following these relationships, we adopt the three-zone approach and say that the forecast model is in the red region if the null hypothesis of sub-calibration is rejected, in the green region if the null hypothesis of super-calibration is rejected, and in the yellow region if none of these hypotheses are rejected. For total ES, we replace $\operatorname{V}_{t}^{\text{VaR}}$ in the above analysis with $\operatorname{V}_{t}^{\text{ES}}=\operatorname{V}^{\text{ES}}\left((\widehat{VaR}_{t},\widehat{ES}_{t}),S_{t}\right)$, where:
@@
This identification function maintains the above relationships; see Section 2.2.2 of Nolde and Ziegel (2017). For $j$th ES contribution, $j\in\{1,\dots,d\}$, we replace $\operatorname{V}_{t}^{\text{VaR}}$ in the above analysis with $\operatorname{V}_{j,t}^{\text{ESC}}=\operatorname{V}_{j}^{\text{ESC}}\left((\widehat{ESC}_{j,t},\widehat{VaR}_{t}),\bm{X}_{t}\right)$, where $\operatorname{V}_{j}^{\text{ESC}}$ is as defined in Proposition 2. We report the results of these tests in Tables 3 and 4. Due to joint identifiability of total ES (see Section 2.1 of  Nolde and Ziegel 2017) and $j$th ES contribution for $j=1,\dots,d$ (see Proposition 2) in combination with total VaR, we also conduct Wald-tests for these risk quantities in the one-step approach; see Section S2.4 for details.
@@
GARCH.BU 0.689 0.620 0.690 0.310 yellow
GARCH.TD -0.073 0.949 0.474 0.526 yellow
@@
  • ${}^{a}$he p-values are calculated based on the three different null hypotheses, where “$=$ True” means that the forecast is precise, $\leq$ True” represents the hypothesis that the forecast is under-estimated, and $\geq$ True” stands for the case when the forecast is over-estimated.
  • ${}^{b}$ Results of the three-zone approach are presented.
Table 4: Results of the DM-type tests to verify the forecast accuracy of the $j$th ES contribution for $j=1,\dots,d$. See the footnotes in Table 3 for details.
@@
GARCH.BU -0.005 0.742 0.371 0.629 yellow
GARCH.TD -0.008 0.565 0.283 0.717 yellow
@@
CR.LSE -0.009 0.541 0.271 0.729 yellow
CR.OPT -0.012 0.363 0.182 0.818 yellow
(2) ES contribution (GOOGL)
@@
GARCH.BU -0.004 0.731 0.366 0.634 yellow
GARCH.TD -0.009 0.427 0.214 0.786 yellow
@@
CR.LSE -0.008 0.474 0.237 0.763 yellow
CR.OPT 0.001 0.938 0.531 0.469 yellow
(3) ES contribution (TSLA)
@@
GARCH.BU -0.063 0.028 0.014 0.986 green
GARCH.TD -0.023 0.365 0.183 0.817 yellow
@@
CR.LSE -0.012 0.636 0.318 0.682 yellow
CR.OPT -0.017 0.458 0.229 0.771 yellow
We finally visually check the performance of the forecasts by Murphy diagrams. For total VaR, the diagram is written based on (4); see Ehm et al. (2016) for details. We refer to Ziegel et al. (2020) for the Murphy diagram of ES. The Murphy diagram of the tuple of ES contributions is drawn based on Proposition 3 (M2). We also display Murphy diagrams of the $j$th ES contribution for $j=1,\dots,d$ based on Proposition 3 (M1). The results appear in Figure 1.
@@
Figure 1: Murphy diagrams of total VaR, total ES, tuple of ES contributions, and each of them with confidence level of $0.975$ estimated by the $6$ models.
@@ 5.3 Discussion
According to the results of the comparative backtests in Tables 1 and 2, we first observe that HS is inferior to others in terms of average scores, which is detected by the red region in the three-zone approach only except for the ES contribution of AMZN. Moreover, CR.OPT tends to have lower scores than others and achieves the best performance for total VaR, tuple of ES contributions, and ES contribution of TSLA. Overall, GARCH.TD outperforms GARCH.BU, and HAR.GARCH does not perform well except for the ES contribution of AMZN. For total VaR in Table 1 (1), we observe the tendency that the forecasts of total VaR used in CR.LSE and CR.OPT are more accurate than others. Nevertheless, no test statistically supports the difference in the forecast accuracy of total VaRs except for HS, which justifies the forecast comparisons of ES contributions in terms of their scores.
@@
In Table 1 (2) for total ES we find slightly lower performances of GARCH.BU and HAR.GARCH than others, which indicate that the top-down approach can be slightly more preferable than the bottom-up approach, and it may not be recommendable to estimate total ES indirectly as the sum of ES contributions. In addition to total ES, both of GARCH.BU and HAR.GARCH do not perform well for the tuple of ES contributions as well as each of them except for AMZN. For the case of AMZN, HAR.GARCH performs the best, which may imply the existence of the hysteretic effect and/or dynamic conditional correlations in the series of stock returns. From the perspective of multi-objective elicitability, we can directly compare the forecast accuracy of ES contributions between CR.OPT and CR.LSE since they share common forecasts of total VaR. We observe that CR.OPT outperforms CR.LSE for all the cases except the ES contribution of GOOGL. Finally, we see in Table S5 of Section S2.4 that the Wald-tests do not statistically support any additional orders among our compared models, which may not be surprising since no difference in the forecast accuracy of total VaRs is statistically verified other than HS.
@@
We next check the results of the traditional backtests in Tables 3 and 4. Overall, the results are consistent with those of the comparative backtests, and HS is particularly not identified well. Other than HS, we find that GARCH.BU and HAR.GARCH are two models that typically fail the tests. For these models, the forecasts are often under-estimated except for the case that GARCH.BU over-estimates the ES contribution of TSLA. The other three models, GARCH.TD, CR.LSE, and CR.OPT, pass the traditional backtests. These results are consistent with those in Table S6 of Section S2.4. The only difference in this one-step approach is that the Wald-test for GARCH.TD rejects the null hypothesis of under-estimation of total ES with a correct specification of total VaR.
@@
Since all the above results depend on the specific choices of scoring functions, it is beneficial to check the Murphy diagrams to diagnose the robustness of our observations for different choices of scoring functions. Overall, we do not observe any uniform dominance among the curves except HS, which is distinguishable for all the risk quantities. Compared with total VaR and total ES, larger differences are more visible for ES contributions. Particularly for the tuple of ES contributions and the ES contribution of TSLA, the curves of CR.OPT are typically lower, and those of GARCH.BU and HAR.GARCH tend to be higher. In addition, the curves are more tangled for ES contributions of AMZN and GOOGL. These differences are consistent with the fact that, from Table 2 and the y-axes of Murphy diagrams of ES contributions, TSLA has much larger contribution to the score of the tuple of ES contributions than AMZN and GOOGL have, and thus optimization in CR.OPT puts higher weight to the score of TSLA under the full allocation property.
@@ 6 Conclusion and outlook
We conduct traditional and comparative backtests for the gradient allocations of ES based on newly developed Murphy diagrams, which are robust diagnostic tools against the choice of scoring function. Moreover, motivated by the comparability of ES contributions in multi-objective elicitability and by the preservability of the full allocation property, we propose a novel semiparametric compositional regression model to estimate the tuple of ES contributions. Our empirical analysis demonstrates superior performance of our proposed model for forecasting ES contributions and the induced total ES. We believe that these results are of great benefit for portfolio and enterprise risk management in financial and regulatory applications.
We conclude this section by offering an outlook for future research on backtesting ES contributions. First, exploring theoretical aspects of our proposed compositional regression model, such as stability, consistency, and asymptotic normality, is of great interest. As a practical aspect, variable selection for our proposed model lies beyond the scope of the present paper. Second, in a future work it may be interesting to extend our estimation procedure of risk allocations to VaR contributions, the gradient allocation of VaR (see Gribkova et al. 2023; Koike et al. 2022, for recent works), and other capital allocation rules based on optimization (Dhaene et al. 2012; Maume-Deschamps et al. 2016; Koike and Hofert 2021).
@@ Acknowledgements
Takaaki Koike was supported by Japan Society for the Promotion of Science (JSPS KAKENHI Grant Number JP21K13275). Cathy W.S. Chen’s research is funded by the National Science and Technology Council, Taiwan (NSTC112-2118-M-035-001-MY3).
@@
Section S1 provides proofs and details of the technical statements in the main paper. Section S2 gives a for detailed description of the empirical analysis in the main paper as well as some additional results of backtesting in the one-step approach (Fissler and Hoga, 2023). Throughout the supplementary material, references starting with “S” refer to this supplement, and those without this prefix refer to the main paper.
Appendix S1 Technical details
S1.1 Proof of Proposition 3
(M2) is an immediate consequence of (M1), and thus it suffices to show (M1). We fix $j\in\{1,\dots,d\}$. For a strictly convex function $\phi_{j}:\mathbb{R}\rightarrow\mathbb{R}$ with subgradient $\phi_{j}^{\prime}$, define the Bregman-type function $\Phi_{j}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ as:
$\Phi_{j}(m,x)=\phi_{j}(x)-\phi_{j}(m)-\phi_{j}^{\prime}(m)(x-m).$
Following the proof of Theorem 1 in Ehm et al. (2016), it holds that:
$=\mathbbm{1}{\{s>v\}}\Phi_{j}(m,x_{j})$
S1.2 Order sensitivity
In this section we explore order sensitivity for ES contributions, which concerns whether two misspecified forecasts of ES contributions are ordered by means of the multi-objective scoring function in Proposition 1 if one tuple of forecasts is closer to the true ES contributions than another. The next proposition presents conditions under which two misspecified forecasts are ordered in this sense.
@@
We classify an approach of building dynamic models of ES contributions into the bottom-up and top-down approaches. In the bottom-up approach we estimate the risk functional of $\bm{X}_{n+1}|\mathcal{G}_{n}$ based on a joint model $\{\bm{X}_{t}\}_{t\in\mathbb{T}}$. This joint model determines the dynamics of $\{S_{t}\}_{t\in\mathbb{T}}$ and $\{(X_{j,t},S_{t})\}_{t\in\mathbb{T}}$, based on which we predict total VaR, total ES, and ES contributions at time $n+1$. An example is the bottom-up GARCH model described in Section S2.2. A major advantage of this approach is that the estimated model can estimate quantities beyond ES contributions. This versatility, however, comes at the expense of possibly low forecast accuracy of total VaR and ES since the aggregate loss is modeled only indirectly. Another potential drawback is the challenge of modeling $d$-dimensional time series especially when $d$ is large.
In the top-down approach we first estimate the dynamics of total VaR on $\mathbb{T}$ and then estimate the dynamics of ES contributions on $\mathbb{T}$ based on some necessary model specification. For example, once the dynamics of total VaR are specified, that of the $j$th ES contribution can be estimated only from the bivariate time series $\{(X_{j,t},S_{t})\}_{t\in\mathbb{T}}$ for each $j\in\{1,\dots,d\}$ without specifying the joint model $\{\bm{X}_{t}\}_{t\in\mathbb{T}}$. Total ES can be predicted together with total VaR in the first stage or estimated as the sum of the ES contributions estimated in the second stage. Compared with the bottom-up approach, the potential benefits of the top-down approach are flexibility of the VaR model and a reduction of modeling effort to estimate ES contributions. On the other hand, this model specification can be involved in the compatibility problem that concerns the existence of a joint model $(\bm{X}_{t},S_{t})$ whose bivariate marginal time series $(X_{j,t},S_{t})$, $j=1,\dots,d$, are the specified ones. Complexity of this compatibility problem may hinder the use of the estimated model other than the forecasting problem of ES contributions.
@@
Historical simulation (HS): In this model the estimator $\widehat{VaR}_{n+1}^{\text{HS}}$ is given as the empirical $\alpha$-quantile based on $S_{1},\dots,S_{n}$. We then predict the total ES and ES contributions as follows:

### v2

@@ 5.2 Model evaluation
(1) ESC (AMZN)
HS 33.603 6 0.091 0.954 0.046 (red)
GARCH.BU 18.646 2 0.857 0.428 0.572 yellow
GARCH.TD 19.951 4 0.784 0.608 0.392 yellow
HAR.GARCH 13.351 1 0.495 0.247 0.753 yellow
CR.LSE 23.295 5 0.152 0.924 0.076 yellow
CR.OPT 19.291 3 —- —- —- —-
(2) ESC (GOOGL)
HS 28.672 6 0.083 0.959 0.041 (red)
GARCH.BU 13.291 4 0.496 0.752 0.248 yellow
GARCH.TD 11.865 3 0.963 0.518 0.482 yellow
HAR.GARCH 15.629 5 0.285 0.857 0.143 yellow
CR.LSE 11.054 1 0.658 0.329 0.671 yellow
CR.OPT 11.782 2 —- —- —- —-
(3) ESC (TSLA)
HS 126.941 6 0.011 0.994 0.006 (red)
GARCH.BU 86.260 4 0.011 0.994 0.006 red
GARCH.TD 62.882 3 0.257 0.871 0.129 yellow
HAR.GARCH 95.799 5 0.130 0.935 0.065 yellow
CR.LSE 62.115 2 0.078 0.961 0.039 red
CR.OPT 54.765 1 —- —- —- —-
We next describe our traditional backtests. Denote by $\left(\widehat{ESC}_{1,t},\dots,\widehat{ESC}_{d,t},\widehat{VaR}_{t},\widehat{ES}_{t}\right)$, $t\in\mathbb{T}_{\text{out}}$, the tuple of forecasted risk functionals. For total VaR, let $\operatorname{V}_{t}^{\text{VaR}}=\operatorname{V}^{\text{VaR}}\left(\widehat{VaR}_{t},S_{t}\right)$, where $\operatorname{V}^{\text{VaR}}$ is as defined in Proposition 2. We construct the DM-type test statistic $\sqrt{T}\bar{\operatorname{V}}_{T}^{\text{VaR}}/\hat{\sigma}_{T}^{\text{VaR}}$, where $\bar{\operatorname{V}}_{T}^{\text{VaR}}=(1/T)\sum_{t=1}^{T}\operatorname{V}_{t}^{\text{VaR}}$, and $\hat{\sigma}_{T}^{\text{VaR}}$ is the HAC estimator of the standard deviation of $\sqrt{T}\bar{\operatorname{V}}_{T}^{\text{VaR}}$ as considered in the comparative backtests above. Following Section 2.2 of Nolde and Ziegel (2017), we use a normal distribution as a null distribution to test unconditional calibration, sub-calibration, and super-calibration, which are associated with precise estimation, over-estimation, and under-estimation, respectively, for this identification function of VaR; see Section S1.1. Following these relationships, we adopt the three-zone approach and say that the forecast model is in the red region if the null hypothesis of sub-calibration is rejected, in the green region if the null hypothesis of super-calibration is rejected, and in the yellow region if none of these hypotheses are rejected. For total ES, we replace $\operatorname{V}_{t}^{\text{VaR}}$ in the above analysis with $\operatorname{V}_{t}^{\text{ES}}=\operatorname{V}^{\text{ES}}\left((\widehat{VaR}_{t},\widehat{ES}_{t}),S_{t}\right)$, where:
@@
This identification function maintains the above relationships on over- and under-estimation; see Section 2.2.2 of Nolde and Ziegel (2017). For $j$th ESC, $j\in\{1,\dots,d\}$, we replace $\operatorname{V}_{t}^{\text{VaR}}$ in the above analysis with
$\operatorname{V}_{j,t}^{\text{ESC}}=\operatorname{V}_{j}^{\text{ESC}}\left((\widehat{ESC}_{j,t},\widehat{VaR}_{t}),\bm{X}_{t}\right),$
where $\operatorname{V}_{j}^{\text{ESC}}$ is as defined in Proposition 2. We report the results of these tests in Tables 4 and 5. Due to joint identifiability of total ES (see Section 2.1 of  Nolde and Ziegel 2017) and $j$th ESC for $j=1,\dots,d$ (see Proposition 2) in combination with total VaR, we also conduct Wald-tests for these risk quantities in the one-step approach; see Section S2.3 for details.
@@
GARCH.BU 0.634 0.645 0.677 0.323 yellow
GARCH.TD 0.508 0.652 0.674 0.326 yellow
@@
  • aThe p-values are calculated based on the three different null hypotheses, where “$=$ True” means that the forecast is precise, $\leq$ True” represents the hypothesis that the forecast is under-estimated, and $\geq$ True” stands for the case when the forecast is over-estimated.
  • b Results of the three-zone approach are presented.
Table 5: Results of the DM-type tests to verify the forecast accuracy of the $j$th ESC for $j=1,\dots,d$. See the footnotes in Table 4 for details.
@@
GARCH.BU -0.005 0.700 0.350 0.650 yellow
GARCH.TD -0.008 0.594 0.297 0.703 yellow
@@
CR.LSE -0.007 0.656 0.328 0.672 yellow
CR.OPT -0.014 0.296 0.148 0.852 yellow
(2) ESC (GOOGL)
@@
GARCH.BU -0.005 0.669 0.335 0.665 yellow
GARCH.TD -0.009 0.393 0.196 0.804 yellow
@@
CR.LSE -0.006 0.550 0.275 0.725 yellow
CR.OPT -0.005 0.643 0.322 0.678 yellow
(3) ESC (TSLA)
@@
GARCH.BU -0.070 0.017 0.009 0.991 green
GARCH.TD -0.023 0.353 0.177 0.823 yellow
@@
CR.LSE -0.015 0.536 0.268 0.732 yellow
CR.OPT -0.009 0.701 0.351 0.649 yellow
In the end, we visually check the performance of the forecasts by Murphy diagrams. For total VaR, the diagram is written based on (4); see Ehm et al. (2016) for details. We refer to Ziegel et al. (2020) for the Murphy diagram of ES. The Murphy diagram of the tuple of ESCs is drawn based on Proposition 3 (M2). We also display Murphy diagrams of the $j$th ESC for $j=1,\dots,d$ based on Proposition 3 (M1). The results appear in Figure 3.
@@
Figure 3: Murphy diagrams of total VaR, total ES, tuple of ESCs, and each of them with confidence level of $0.975$ estimated by the six models.
@@ 5.3 Discussion
According to the results of the comparative backtests in Tables 2 and 3, we first observe that HS is inferior to others in terms of average scores, which is detected by the red region in the three-zone approach. Moreover, CR.OPT tends to have lower scores than others and achieves the best performance for the tuple of ESCs and the ESC of TSLA. Overall, GARCH.TD outperforms GARCH.BU, and HAR.GARCH does not perform well except for the ESC of AMZN. For total VaR in Table 2 (1), we observe the tendency that the forecasts of total VaR used in GARCH.TD, CR.LSE and CR.OPT are more accurate than others. Nevertheless, no test statistically supports the difference in the forecast accuracy of total VaRs except for HS, which justifies the forecast comparisons of ESCs in terms of their scores.
@@
In Table 2 (2) for total ES, we find slightly lower performances of GARCH.BU and HAR.GARCH than others. This indicates that the top-down approach can be slightly more preferable than the bottom-up approach, and it may not be recommendable to estimate total ES indirectly as the sum of ESCs. In addition to total ES, both of GARCH.BU and HAR.GARCH do not perform well for the tuple of ESCs as well as each of them except for AMZN. For the case of AMZN, HAR.GARCH performs the best, which may imply the existence of the hysteretic effect and/or dynamic conditional correlations in the series of stock returns. From the perspective of multi-objective elicitability, we can directly compare the forecast accuracy of ESCs among GARCH.TD, CR.LSE and CR.OPT since they share common forecasts of total VaR. We observe that CR.OPT outperforms CR.LSE for all the cases except the ESC of GOOGL. Finally, we see in Table S6 of Section S2.3 that the Wald-tests statistically support the superiority of CR.OPT to HS for AMZN and GOOGL and to HS and GARCH.BU for the tuple of ESCs and TSLA. These orders are also detected in Table 2. In addition, the Wald-tests do not support the order between CR.LSE and CR.OPT, which is supported in the DM-tests in Tables 2 and 3.
@@
We next check the results of the traditional backtests in Tables 4 and 5. Overall, the results are consistent with those of the comparative backtests, and HS particularly does not identify the true risk quantities well. Other than HS, we find that GARCH.BU and HAR.GARCH are two models that typically fail the tests. For these models, the forecasts are often under-estimated except for the case that GARCH.BU over-estimates the ESC of TSLA. The other three models, GARCH.TD, CR.LSE, and CR.OPT, pass the traditional backtests. These results are consistent with those in Table S7 of Section S2.3, where HS, GARCH.BU and HAR.GARCH typically fail the model validation tests.
@@
Since all the above results depend on the specific choices of scoring functions, it is beneficial to check the Murphy diagrams to diagnose the robustness of our observations for different choices of scoring functions. Overall, we do not observe clear uniform dominance among the curves except HS, which is distinguishable for all the risk quantities. Compared with total VaR and total ES, larger differences are more visible for ESCs. Particularly for the tuple of ESCs and the ESC of TSLA, the curves of CR.OPT are typically lower, and those of GARCH.BU and HAR.GARCH tend to be higher. In addition, the curves are more tangled for ESCs of AMZN and GOOGL. These differences are consistent with the fact that, from Table 3 and the y-axes of Murphy diagrams of ESCs, TSLA has larger contribution to the score of the tuple of ESCs than AMZN and GOOGL have, and thus optimization in CR.OPT should put higher weight to the score of TSLA under FAP. In summary, the presented diagrams visually indicate the superiority of CR.OPT to other models for the tuple of ESCs and the ESC of TSLA. For the ESC of AMZN, and possibly of GOOGL, ranking of models may vary depending on the choice of the scoring function.
@@ 6 Conclusion and outlook
We conduct traditional and comparative backtests for the gradient allocations of ES, and visually check the robustness of our observations based on Murphy diagrams. Moreover, motivated by the comparability of ESCs in multi-objective elicitability and by the preservability of FAP, we propose a novel semiparametric compositional regression model to estimate the tuple of ESCs. Our empirical analysis demonstrates superior performance of our proposed model for forecasting ESCs. We believe that these results are of great benefit for portfolio and enterprise risk management in financial and regulatory applications.
We conclude this section by offering an outlook for future research on backtesting ESCs and our proposed model. First, exploring theoretical aspects of our proposed compositional regression model, such as stability, consistency, and asymptotic normality, is of great interest. As a practical aspect, variable selection for our proposed model and the choice of the transform between the simplex and the real space are beyond the scope of the present paper. The cases of high-dimensional portfolios and of negative or zero ESCs may require further analyses. Moreover, in a future work it may be interesting to extend our estimation procedure of risk allocations to VaR contributions, the gradient allocation of VaR (see Koike et al. 2022; Gribkova et al. 2023, for recent works), and other capital allocation rules based on optimization (Dhaene et al. 2012; Maume-Deschamps et al. 2016; Koike and Hofert 2021). Finally, other backtesting procedures (Banulescu-Radu et al. 2021; Wang et al. 2022; Hoga and Demetrescu 2023; Hué et al. 2024) could bring new insights on backtesting capital allocations.
@@ Acknowledgements
Takaaki Koike was supported by Japan Society for the Promotion of Science (JSPS KAKENHI Grant Numbers JP21K13275 and JP24K00273). Cathy W.S. Chen’s research is funded by the National Science and Technology Council, Taiwan (NSTC112-2118-M-035-001-MY3). Edward M.H. Lin’s research is funded by the National Science and Technology Council, Taiwan (NSTC111-2118-M-029-003-MY2). The authors also thank Tim J. Boonen for constructive comments and suggestions.
Declaration of interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@
Section S1 provides proofs and details of some statements presented in the main paper. Section S2 gives a detailed description of the empirical analysis in Section 5 and some additional results of backtesting ESCs in the one-step approach (Fissler and Hoga 2024). Throughout the supplementary material, references starting with “S” refer to this supplement, and those without this prefix refer to the main paper.
Appendix S1 Proofs and details
S1.1 Order sensitivity and orientation
A key property of scoring functions regarding the comparative backtests is order sensitivity, which concerns whether two misspecified forecasts of ESCs are ordered by means of the multi-objective scoring function in Proposition 1 if one tuple of forecasts is closer to the true ESCs than another. The next proposition presents conditions under which two misspecified forecasts are ordered in this sense.
@@
To summarize it briefly, the average scores of the two misspecified forecasts of ESCs are ordered when (i) the corresponding total VaRs are ordered, or (ii) the misspecified total VaRs are equal (in this case, $\operatorname{S}_{1}^{\text{ESC}},\dots,\operatorname{S}_{d}^{\text{ESC}}$ elicit biased ESCs), the two misspecified ESCs are ordered in the component-wise sense, and they are not between the true and biased ESCs. Regarding the latter case, misspecification of total VaR produces intervals such that forecasts of ESCs belonging there have inconsistent orders of average scores.
@@
Regarding the concept of orientation (Steinwart et al. 2014), it is useful to associate the sign of $\mathbf{\bar{V}}$ with over- and under-estimations of ESCs since under-estimation of risk functionals is considered to be more problematic than over-estimation from a regulatory viewpoint. With the identification functions in Proposition 2, we have, for $\bm{X}\in\mathcal{F}$, that $\mathbb{E}[\operatorname{V}^{\text{VaR}}(v,S)]\leq 0$ if and only if $\operatorname{VaR}_{\alpha}(S)\leq v$; i.e., the forecast $v$ over-estimates the total VaR. Moreover, under the correct specification $v=\operatorname{VaR}_{\alpha}(S)$ and for $j\in\{1,\dots,d\}$, a forecast $m_{j}\in\mathbb{R}$ of the $j$th ESC is over-estimated; i.e., $\operatorname{ESC}_{\alpha}(X_{j},S)\leq m_{j}$, if and only if $\mathbb{E}[\operatorname{V}_{j}^{\text{ESC}}((m_{j},v),\bm{X})]\leq 0$.
S1.2 Proof of Proposition 3
(M2) is an immediate consequence of (M1), and thus it suffices to show (M1). We fix $j\in\{1,\dots,d\}$. For a strictly convex function $\phi_{j}:\mathbb{R}\rightarrow\mathbb{R}$ with derivative $\phi_{j}^{\prime}$, define the Bregman-type function $\Phi_{j}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ as:
$\Phi_{j}(m,x)=\phi_{j}(x)-\phi_{j}(m)-\phi_{j}^{\prime}(m)(x-m).$
Following the proof of Theorem 1 in Ehm et al. (2016), it holds that:
$=\mathbbm{1}{\{s>v\}}\Phi_{j}(m,x_{j})$
S1.3 Interpretation of the elementary scoring function
The elementary scoring function (6) can be interpreted as a degree of regret in the following setting. Suppose that the $j$th branch of a company has a fixed capital $\eta$ to cover a future loss, whose point forecast is denoted by $m_{j}$, incurred at this branch under the specific situation when the company incurs a loss greater than $v$. If $m_{j}\leq\eta$, then the branch expects that the initial capital can cover a future loss. In this case, since an excess loss $(x_{j}-\eta)$ is incurred when the realized loss $x_{j}$ exceeds $\eta$ and $s>v$ occurs, the amount $\operatorname{S}_{j,\eta}^{\text{ESC}}((m_{j},v),\bm{x})=\mathbbm{1}{\{s>v\}}(x_{j}-\eta)_{+}$ can be understood as a degree of regret against the initial expectation. If $\eta<m_{j}$, then the branch expects that the initial capital does not cover a future loss, and thus some risk treatment can be conducted. In this case, the branch realizes the squandered opportunity of capital reduction $(\eta-x_{j})_{+}$ when $s>v$ occurs, and thus the degree of regret against the initial expectation can be measured by $\operatorname{S}_{j,\eta}^{\text{ESC}}((m_{j},v),\bm{x})=\mathbbm{1}{\{s>v\}}(\eta-x_{j})_{+}$. Provided $v=\operatorname{VaR}_{\alpha}(S)$, the true $j$th ESC is the optimal $m_{j}$ minimizing the expected degree of regret.
@@
We classify an approach of building dynamic models of ESCs into the bottom-up and top-down approaches. In the bottom-up approach we estimate the risk functional of $\bm{X}_{n+1}|\mathcal{G}_{n}$ based on a joint model $\{\bm{X}_{t}\}_{t\in\mathbb{T}}$. This joint model determines the dynamics of $\{S_{t}\}_{t\in\mathbb{T}}$ and $\{(X_{j,t},S_{t})\}_{t\in\mathbb{T}}$, based on which we predict total VaR, total ES, and ESCs at time $n+1$. An example is the bottom-up GARCH model described in Section S2.2. A major advantage of this approach is that the estimated model can be used to estimate quantities other than ESCs. This versatility, however, comes at the expense of possibly low forecast accuracy of total VaR and ES since the aggregate loss is modeled only indirectly. Another potential drawback is the challenge of modeling $d$-dimensional time series especially when $d$ is large.
In the top-down approach we first estimate the dynamics of total VaR on $\mathbb{T}$ and then estimate the dynamics of ESCs on $\mathbb{T}$ based on some necessary model specification. For example, once the dynamics of total VaR is specified, that of the $j$th ESC can be estimated only from the bivariate time series $\{(X_{j,t},S_{t})\}_{t\in\mathbb{T}}$ for each $j\in\{1,\dots,d\}$ without specifying the joint model $\{\bm{X}_{t}\}_{t\in\mathbb{T}}$. Total ES can be predicted together with total VaR in the first stage or estimated as the sum of the ESCs estimated in the second stage. Compared with the bottom-up approach, the potential benefits of the top-down approach are flexibility of the VaR model and reduction of modeling effort to estimate ESCs. On the other hand, this model specification can be involved in the compatibility problem that concerns the existence of a joint model $(\bm{X}_{t},S_{t})$ whose bivariate marginal time series $(X_{j,t},S_{t})$, $j=1,\dots,d$, are the specified ones. Complexity of this compatibility problem may hinder the use of the estimated model other than the forecasting problem of ESCs.
@@
Historical simulation (HS): In this model the estimator $\widehat{VaR}_{n+1}^{\text{HS}}$ is given as the empirical $\alpha$-quantile based on $S_{1},\dots,S_{n}$. We then predict the total ES and ESCs as follows: